OpenStudy (anonymous):

Please help. studying for my exam and not sure how to do this on my calculator. solve the inequality ln x ≤.8^x using your calculator. i've been putting in ln(x)-.8^x in my calculator. i know the answer is (0,1.9188] but not sure how to get it

OpenStudy (klimenkov):

You have an inequality\[\ln x\le8^x\]and want to solve it?

OpenStudy (anonymous):

yes, I'm supposed to do it on a calculator but I'm not sure how. I have a ti-84

OpenStudy (klimenkov):

Its a pity, but i dont know the abilities of your calculator. Can it solve any inequalities?

OpenStudy (anonymous):

well i know that i put it into the calculator and graph it but from there I'm not sure how to get (0, 1.9188] cause that is the answer we were given by the teacher.

OpenStudy (klimenkov):

Is it possible that your teacher has made a mistake?

OpenStudy (anonymous):

no i know that there isa way to do it on the calculator I'm just not very good with it. thanks for trying to help though

OpenStudy (klimenkov):

Hm. Your inequality says that you have to find all the x, where lnx<=8^x. But if you plot 2 graphs you will see that this graphs dont intersect. And lnx is always lower than 8^x. So your inequality is true for all x from [0, infinity]

OpenStudy (klimenkov):

Oh, im really sorry. It must be 0.8 instead of 8. So, now I ask you can your calculator solve any equations?

OpenStudy (anonymous):

yes its .8, and I'm pretty sure my calculator can solve equations

OpenStudy (klimenkov):

1. Graph both of lnx and 0.8^x on your calculator. 2. Notice that the graph of ln x is lower than the graph of 8^x when x is from zero to the point of intersection. 3. Find the point of intersection. Put in your calculator equation ln x = 8^x. Get the solution. 4. So the answer is (0, the point of intersection]. We include this point because inequality is unstrict.

OpenStudy (anonymous):

I'm still not getting it on my calc. thank you for all your help though

OpenStudy (klimenkov):

Where do you have a problem?

OpenStudy (anonymous):

i can't get the intersect

OpenStudy (klimenkov):

Can you solve an equation\[\ln x =0.8^x\]using your calculator?